# Human Brain Exhibits Distinct Patterns When Listening to Fake Versus Real Audio: Preliminary Evidence

### v2

@@ 1 Introduction
Despite the many remarkable successes of machine learning (ML), there is still a gap between human and machine learning. Evidence for this gap is that patterns of errors made by today’s machine learning algorithms are different from the patterns of humans performing the same tasks. While the human brain has been a source of inspiration for machine learning, little effort has been made to directly use data collected from brain activities as a guide for machine learning algorithms fong2018 . To address the generalisation problem outlined above, this paper proposes anomaly detectors (i.e., cloned voice detectors) that are guided by human brain activities (i.e., EEG). The hypothesis is that the neural activity carries information for the concept of anomaly (i.e., “meta level”) which is agnostic to the exact type of the anomaly (i.e., “surface level”). This hypothesis is supported by a recent study suggesting that we can reliably decode deepfake content using human neural activity moshel2022you .
@@ 2.1 Problem Formulation
Given a labeled audio time series dataset $A$ and an equivalent labeled EEG time series dataset $E$ of a human subject listening to $A$, the problem is how …
@@ 2.2 Data Collection
Real Audio Data
We collected audio data with ethics clearance for deepfake audio production. We hired 20 native English-speaking actors (10 males, 10 females) between the age of 20-35. We prepared 10 conversational scripts with controlled topics for producing 10 single-turn (i.e., monologue) audios. Each script roughly took 3 minutes to read. We required to collect the data ourselves (as opposed to using publicly available alternatives) to have full control over topics, prosodic features and have the ethical consents to release this unique dataset paired with its equivalent audio (see next paragraph) to the research community.
@@
The total duration of our mixed audio file is 44 minutes and 26 seconds including the pauses -silent segments (refereed to as “silence audio”) that were added before and after each script within the audio file. The duration of the silence audio is 1 minute and the reason why we added them is to give the human participants an opportunity to rest during the EEG data collection phase.
@@ 3 Experimental Results
We visualize the Highest Density Regions (HDR;hyndman1996computing ) of the samples of EEG time series ($E_{i}$s), as well the HDR of the representations of audios from the classification head of the CITE state-of-the-art audio deepfake detection system using the Mapper [^2] https://github.com/scikit-tda/kepler-mapper algorithm singh2007topological . The HDR cuts the overall density space to create latent spaces that contain above a threshold probability mass (i.e., minimum ditance $\geq 0.65$ ). The Mapper’s output is represented as a graph in which each component corresponds to a cluster of points that together form a high density region. The connectivity of the graph reflects some topoligical characteristics of the representation space, with darker colors (i.e., dark purples reflecting higher density). Blue represents fake and purple represents real. Fig. 4 demonstrates the Mapper output on the audio representations versus EEG representations. The observation highlights a much more clear separation of fake and real signals in the EEG space compared to the audio’s. This is an strong indication that EEG’s representation of deepfake audio could offer a more discriminative representation of data.

### v3

@@ 1 Introduction
Despite the many remarkable successes of machine learning, there is still a gap between human and machine learning. Evidence for this gap is that patterns of errors made by today’s machine learning algorithms are different from the patterns of humans performing the same tasks. While the human brain has been a source of inspiration for machine learning, little effort has been made to directly use data collected from brain activities as a guide for machine learning algorithms fong2018 . To address the generalisation problem outlined above, this paper proposes anomaly detectors (i.e., cloned voice detectors) that are guided by human brain activities (i.e., EEG). The hypothesis is that the neural activity carries information for the concept of anomaly (i.e., “meta level”) which is agnostic to the exact type of the anomaly (i.e., “surface level”). This hypothesis is supported by a recent study suggesting that we can reliably decode deepfake content using human neural activity moshel2022you .
@@ 2.1 Problem Formulation
Given a labeled audio time series dataset $A$, and an equivalent labeled EEG time series dataset $E$ of a human subject listening to $A$, where a label of each sentence in $A$ belong to a set $labels=\{real,fake\}$, the problem is to train a classifier on $E$ to classify $A$ using $labels$. We formulate this as a time series classification problem.
@@ 2.2 Data Collection
Real Audio Data.
We collected audio data with ethics clearance for deepfake audio production. We hired 20 native English-speaking actors (10 males, 10 females) between the age of 20-35. We prepared 10 conversational scripts with controlled topics for producing 10 single-turn (i.e., monologue) audios. Each script roughly took 3 minutes to read. We required to collect the data ourselves (as opposed to using publicly available alternatives) to have full control over topics, prosodic features and have the ethical consents to release this unique dataset to the research community.
@@
To generate high-quality, realistic fake audio, we employed two state-of-the-art speech generation methods, VITS kimConditional2021 and YourTTS casanovaYourTTS2022 . VITS is a few-shot speech generation method capable of producing very realistic fake voice through fine-tuning on real data. After fine-tuning, the speaker-specific VITS models receive text inputs and generate fake audio based on the styles of the target speaker. YourTTS is a zero-shot speech generation method that can mimic any speaker’s style without fine-tuning. The YourTTS model is pre-trained on multiple datasets to ensure generalizability. The YourTTS model takes text and reference audio as input and generates fake audio with a similar style to the reference audio.
To analyze the impact of different quality settings of the deepfake voices, we defined three quality configurations:
  • High: Fine-tuning VITS on all recorded real speech, then generating target audio included in the training set.
  • Medium: Fine-tuning VITS on the first minute of recorded real speech, then generating target audio that might not be included in the training set.
  • Low: Generating fake audio using YourTTS casanovaYourTTS2022 without fine-tuning.
Inspired by caiYou2022 ; CAI2023103818 ; zhangInitial2021 , we replaced a real segment in the real audio with a fake segment to analyze how humans respond to deepfake audio. We randomly chose the position for insertion from three options: 1) After one minute of real audio; 2) In the middle of the second minute of real audio; and 3) At the end of the real audio. Additionally, we only considered replacing segments with more than five words. Subsequently, we applied the speech generation methods to generate three corresponding fake audios of different quality. In the final step, following caiYou2022 , we normalized the loudness of fake and real segments to ensure smooth transitions between them. Finally, we concatenated all segments to build the final fake audio samples.
@@ 2.3 Audio Preprocessing
Given that the raw audio recordings were noisy, we performed pre-processing to normalize the audio. Firstly, we manually edited the raw audio to remove repetitions, mistakes, and unrelated speech, ensuring the content was clean for input into the speech generation system. Subsequently, we partitioned the audio into several segments based on the transcripts to create clean audio-text pairs for each speaker.
@@ 3 Experimental Results
We visualize the Highest Density Regions (HDR;hyndman1996computing ) of the samples of EEG time series ($E_{i}$s), as well the HDR of the representations of audios from the classification head of the state-of-the-art audio deepfake detection system [^2] https://huggingface.co/HyperMoon/wav2vec2-base-960h-finetuned-deepfake using the Mapper [^3] https://github.com/scikit-tda/kepler-mapper algorithm singh2007topological . The HDR cuts the overall density space to create latent spaces that contain above a threshold probability mass (i.e., minimum ditance $\geq 0.65$ ). The Mapper’s output is represented as a graph in which each component corresponds to a cluster of points that together form a high density region. The connectivity of the graph reflects some topoligical characteristics of the representation space, with darker colors (i.e., dark purples reflecting higher density). Blue represents fake and purple represents real. Fig. 4 demonstrates the Mapper output on the audio representations versus EEG representations. The observation highlights a much more clear separation of fake and real signals in the EEG space compared to the audio’s. This is an strong indication that EEG’s representation of deepfake audio could offer a more discriminative representation of data.